OpenStudy (anonymous):

By using the table in the handbook, the present value of $12,000 for six years compounded at 6 percent semiannually is

OpenStudy (campbell_st):

semi annually for six years will give 12 time periods. use the formula \[A = P(1 + \frac{r}{100})^n\] A = future value P = Principal P = 12,000 r = interest rate r = 6 n = time periods n = 12 substitute and then find th value

OpenStudy (anonymous):

A = 12,000 ( 1 + .06/2) ^12